OpenStudy (anonymous):

An ore car of mass 36000 kg starts from rest and rolls downhill on tracks from a mine. At the end of the tracks, 21 m lower vertically, is a horizontally situated spring with constant 5.1 × 105 N/m. The acceleration of gravity is 9.8 m/s 2 . Ignore friction. How much is the spring compressed in stopping the ore car? Answer in units of m

OpenStudy (anonymous):

I think the 2 steps that you have to take are find the force so that you find the extension. For force : Force or Weight =mg= 36000*9.8= 352800N For compression K=F/x, where F is weight, x is extension and K is contant. x=F/K= 352800/5.1*10^5=

OpenStudy (anonymous):

0.69m